# thanks: Talento-CM fellow

### v1

@@ Abstract
Constraining cosmological parameters through galaxy clusters requires an accurate theory that involves nonlinear scales. Press-Schechter (PS) theory[23] firstly provides a universal formulation for describing the HMF Eq.(2): $f(\nu)=\sqrt{\frac{2}{\pi}}\nu e^{-\nu^{2}/2}$. Here, $\nu\equiv\delta_{\text{crit}}/\sigma(M,z)$ is peak-height, a measure of the rarity of halo, and $\delta_{\text{crit}}=1.686$ is the critical density for halo collapse, weakly dependent on the underlying cosmological model. Universality means that cosmological and redshift dependence enter the equation only through the peak height $\nu$, without affecting functional form [23, 24, 42]. However, subsequent N-body simulations have revealed deviations from universality in redshift [26, 43, 44] and in cosmological model [44].
@@
Regarding cosmological models incorporating massive neutrinos, [45, 46, 47] have discovered that the impact of neutrinos on universality can be minimized by replacing the total matter density $\rho_{m}=\rho_{cdm+b+\nu}$ in Eq.(2) and Eq.(5) with the cold dark matter + baryon density $\rho_{cdm+b}$, [48, 49, 50], and replacing the total matter power spectrum $P_{m}=P_{cdm+b+\nu}$ in Eq.(3) with the cold dark matter + baryon power spectrum $P_{cdm+b}$. The physical reason behind this is that the clustering of neutrinos around the CDM overdensity is minimal and is unlikely to impact nonlinear collapse [51, 52].
By employing this CDM prescription, [47] shows that for simulations with massive neutrinos the HMF is well reproduced by the fitting formula of Tinker et al. [26], which we utilize in this paper.
@@
According to the peak-background split (PBS) approach [53], the bias between the halo and the matter overdensities $b(M,z)$ in Eq.(17) can be derived from the HMF. Therefore, the prescription remains the same, i.e. replacing $P_{m}$ in Eq.(3) with $P_{cdm+b}$, which will recover the universality of bias on large scales [45, 46]. We follow [45] and continue to use the fitting formula proposed by Tinker et al. [27].
@@
However, this prescription only partially recovers the universality of bias. The inclusion of massive neutrinos introduces scale dependence in bias due to two effects, both of which have been observed in N-body simulations: (i) the fact that halos trace CDM fluctuations rather than total (CDM + neutrino) matter density fluctuations [45, 46] and (ii) the scale-dependent growth of CDM density perturbations [54, 55] The former can also be eliminated by the CDM prescription, i.e. replacing $P_{m}$ in Eq.(II.2) with $P_{cdm+b}$. While, the latter, called growth-induced scale-dependent bias (GISDB) effect [56] requires further consideration. To incorporate the GISDB effect, we employ a fitting formula from Eq.(49) of [57], which is based on the spherical collapse model and the PBS approach [58, 57]. This method has been shown to be consistent with the N-body simulation [54]. Importantly, the scale-dependent bias obtained in [57] is nearly independent of mass or redshift:
@@
where $\Delta_{\Lambda CDM}=4.8\times 10^{-3}$,$\alpha=4$, $\Delta_{q}=1.6$, $\Delta_{L}=0.55f_{\nu}$ with $f_{\nu}\equiv\Omega_{\nu}/\Omega_{m}$ representing the neutrino fraction , $q=5k/k_{\text{fs}}$ related to the free streaming scale Eq.(1). Here, $k_{\text{eq}}=0.015{{\,h\,{\rm Mpc}^{-1}}}$ represents the scale of matter-radiation equality, $k_{\text{ref}}$ is a reference wave number on larger scales. The superscript $L$ denotes Lagrangian bias. Since we use $P_{cdm+b}$ to calculate the cluster power spectrum, the Eulerian bias is given by $b=1+b^{L}$. Finally, we can transform the scale-independent effective bias $b_{\text{eff}}(z)=b_{\text{eff}}(k_{\text{ref}},z)$ from Eq.(17) into a scale-dependent effective bias:
@@
Mistakenly ignoring these two neutrino-induced scale dependences in the bias, will have an impact on cosmological parameter inference in galaxy survey [59, 60, 61, 56]. In this paper, we focus on the latter effect, i.e. GISDB effect in cluster survey.
@@
We employ the Fisher matrix to assess the constraint capabilities of clusters. It is defined as the second derivatives of the logarithm of the likelihood function with respect to the parameters $(p_{\mu},p_{\nu})$:
@@
where $\mathcal{L}$ represents the likelihood function. Crudely speaking, the inverse of the Fisher matrix, denoted as $\mathcal{F}^{-1}$, can be considered as the best possible covariance matrix for the measurement errors on the parameters.
Assuming the observed cluster number count $D$ as a Poisson distribution with an expectation and variance $N$ [62], the logarithm of the likelihood function $\mathcal{L}(\bm{D})$ can be expressed as:
@@
where the subscripts $\{i,j\}$ represent two-dimensional $k^{\text{fid}}$-space cells and observed redshift bins, respectively.
@@
Table 1: Parameters and fiducial values.
@@
$\Omega_{m}h^{2}$ Matter density 111including massive neutrinos 0.14240
@@
$A_{\lambda}$ Amplitude in $\ln\lambda^{ob}\right>$ 222mass-richness relation Eq.(II.1) 79.8
$B_{\lambda}$ Mass dependence in $\ln\lambda^{ob}\right>$ 2 0.93
$C_{\lambda}$ Redshift dependence in $\ln\lambda^{ob}\right>$ 2 -0.49
$D_{\lambda}$ Intrinsic scatter in $\sigma_{\ln\lambda^{ob}}$ 333mass-richness relation Eq.(13) 0.217
$\kappa$ Projection effect term in $\sigma_{\ln M^{ob}}$ 444projection effect Eq.(14) 0.01
$B_{M0}$ Constant term in $M_{\text{bias}}$ 555mass bias Eq.(11) 0
$\alpha$ Redshift dependence in $M_{\text{bias}}$ 5 0
@@
The basic cosmological model we consider is an extension of the spatially flat $\Lambda$CDM model by adding one massive neutrino, label as $\nu\Lambda$CDM model. We also consider $\nu w_{0}w_{a}$ model, incorporating dynamic dark energy (DE) with an equation of state (EoS) represented by $w(a)=w_{0}+(1-a)w_{a}$ [65, 66]. The set of cosmological parameters encompasses:
@@
which denotes the baryon density, matter density, neutrino density, DE density, DE Eos parameters, scalar power-law index and present-day normalization of the primordial power spectrum, respectively. The fiducial values for $\{\Omega_{b}h^{2},\Omega_{m}h^{2},\Omega_{DE},n_{s},\sigma_{8}\}$ are following [67], while $\Omega_{\nu}h^{2}=0.0006442$ corresponding to $M_{\nu}=0.06$ eV through:
@@
Additionally, there exists a set of nuisance parameters associated with the scaling relation (SR):
@@
where the first four parameters $\{A_{\lambda},B_{\lambda},C_{\lambda},D_{\lambda}\}$ pertain to the mass-richness relation in Eq.(II.1) and Eq.(13), and their fiducial vales follow Table 2 of [36]: { 79.8, 0.93, -0.49, 0.217}. The projection effect term in Eq.(14) $\kappa=0.1^{2}$ due to a possible contamination rate of about 10% at low redshift for massive clusters for CSST [30, 68]. The mass bias parameters in Eq.(11) are specified as $B_{M0}=0$ and $\alpha=0$.
@@
For CSST, the solid angle is $\Delta\Omega=5.33$, corresponding to the wide survey area of $17500$ deg2. The cluster redshift uncertainty in Eq.(8) is assumed to be $\sigma_{z}=0.001$, considering that the photometric redshift uncertainty for most galaxies is expected to be around $0.02$ for CSST. The incorporation of additional slitless grating spectroscopic data will leads to a smaller one, and the redshift uncertainty of each cluster as a whole would be reduced after averaging the redshifts of the constituent galaxies. Thus, we adopt an optimistic value of $0.001$. The observed redshift is assumed to be up to $z_{\max}=1.5$, using either redshift-based method [69, 70] or color-based method [71]. The observed mass threshold is $M_{200m}\geq 0.836\times 10^{14}{{\,h^{-1}{\rm{M_{\odot}}}}}$, converted from $M_{500c}\geq 0.7\times 10^{14}\,\rm{M_{\odot}}$ to guarantee a completeness of $\geq 90\%$ and a purity of $\geq 90\%$ for the CSST cluster sample [69].
The fiducial wave number ranges span from $k^{\text{fid}}_{\perp,\|}=0.005$ to $0.15$  Mpc-1, with a bin size of $\Delta k^{\text{fid}}_{\perp,\|}=0.005$  Mpc-1. The observed mass bin is $\Delta\ln M^{ob}[\,\rm{M_{\odot}}]=0.2$. For cluster number counts, the observed redshift bin size is $\Delta z^{ob}=0.05$, while for the cluster power spectrum, it is $\Delta z^{ob}=0.2$, ensuring an adequate number of clusters within each redshift bin.
By summing over the aforementioned bins, we obtain the Fisher matrices $\mathcal{F}_{\mu\nu}^{N}$ for the cluster number counts Eq.(27) and $\mathcal{F}_{\mu\nu}^{P}$ for power spectrum Eq.(28). Finally, we can calculate the $1\,\sigma$ error for a parameter $p_{\mu}$ as $\sigma(p_{\mu})=\sqrt{(\mathcal{F}^{-1})_{\mu\mu}}$, as shown in Table 2.
@@
Table 2: Forecasted $1\,\sigma$ errors for all parameters described in subsection II.5. When a parameter is kept fixed, we use ’–’. The header labels the different cosmological models ($\nu\Lambda$CDM and $\nu w_{0}w_{a}$), scenarios (self-calibration and known SR), and probes (cluster number counts $N$, cluster power spectrum $P$, and their combination $N+P$). For complete descriptions, please refer to the main text.
@@
$A_{\lambda}$ 1.5e+02 1.1e+07 1.1e+01 - - - 2.0e+02 1.1e+07 1.2e+01 - - -
$B_{\lambda}$ 1.024 1.7e+05 0.203 - - - 1.356 1.7e+05 0.207 - - -
$C_{\lambda}$ 1.628 1.1e+05 0.328 - - - 2.294 1.0e+05 0.336 - - -
$D_{\lambda}$ 0.363 7.1e+04 0.074 - - - 0.415 7.0e+04 0.075 - - -
$\kappa$ 0.466 5.4e+03 0.129 - - - 0.539 5.3e+03 0.136 - - -
$B_{M0}$ 0.860 6.4e+01 0.029 - - - 1.006 6.3e+01 0.033 - - -
$\alpha$ 0.431 1.9e+03 0.045 - - - 0.434 1.9e+03 0.047 - - -
@@
In this section, we present the main results based on the CSST survey. The redshift distribution of CSST clusters is shown in Figure 1, with a total number count of 408,037. The predicted $1\,\sigma$ errors for all parameters, especially for the total neutrino mass $M_{\nu}$, are shown in Table 2, for the various probes, cosmological models, and scenarios. Specifically, these constraints is derived from the cluster number counts $N$, the cluster power spectrum $P$, and their combination $N+P$, in the two cosmological models: the $\nu\Lambda$CDM model and the $\nu w_{0}w_{a}$ model. To explore the impact of the main systematics in cluster cosmology, i.e. the uncertainty in the SR, we consider two scenarios: (1) the self-calibration scenario, where the SR parameters are jointly constrained with the cosmological parameters; and (2) the known SR scenario, where the SR parameters are fixed.
To express clearly, the constraints on $M_{\nu}$ obtained from $N$, $P$ and $N+P$ are denoted as $\sigma^{N}(M_{\nu})$, $\sigma^{P}(M_{\nu})$ and $\sigma^{N+P}(M_{\nu})$ respectively.
@@ III.1 $\nu\Lambda$CDM model
Figure 2: Constraints on cosmological parameters in the $\nu\Lambda$CDM model, with marginalizing over the SR parameters if exit. All contours denote 68% C.L. Different colors (red, blue and black) represent different probes: the cluster number counts $N$, cluster power spectrum $P$, and their combination $N+P$. Solid and dotted lines represent self-calibration and known SR scenario respectively.
We begin with the $\nu\Lambda$CDM model, referring to columns 2-7 in Table 2. The corresponding contour plot is shown in Fig. 2 for the cosmological parameters $\{M_{\nu},\Omega_{b}h^{2},\Omega_{m}h^{2},\Omega_{DE},n_{s},\sigma_{8}\}$.
It is evident that constraints on all cosmological parameters primarily arise from $P$, except for $\Omega_{DE}$ in the known SR scenario. This is because $N$ is only sensitive to the integration of the matter power spectrum, rather than its shape, resulting in weak constraints on $\Omega_{b}h^{2}$, $n_{s}$ and $M_{\nu}$ [40, 21, 72]. Consequently, the constraint on $M_{\nu}$ from $N$ primarily stems from the influence of neutrinos on the amplitude of the matter power spectrum, leading to a strong degeneracy between $M_{\nu}$ and $\sigma_{8}$. In summary, the constraint on $M_{\nu}$ from $N$ is very weak, only $\sigma^{N}(M_{\nu})=5.215\ (0.706)$ eV in the self-calibration (known SR) scenario.
The power spectrum $P$ serves as a more sensitive probe of the neutrino mass compared to $N$, providing a constraint of $\sigma^{P}(M_{\nu})=0.201\ (0.706)$ eV on its own. When combined with $N$, the improvement factor in the self-calibration (known SR) scenario is 1.72 (6.19).
The incorporation of SR information has a substantial impact on $N$, resulting in an enhancement of constraints on all cosmological parameters by a factor of at least four. Particularly, the constraint on $M_{\nu}$ experiences a significant strength by a factor of 7.4, shifting from $5.215$ to $0.706$ eV. In contrast, the influence of SR information on $P$ is minimal, which is reflected in large errors of SR parameters in the third column, since the SR parameters only affects $b_{\text{eff}}$. Consequently, the enhance in the constraint on $M_{\nu}$ is only 1.04 times, from $0.201$ to $0.194$ eV, when incorporating SR information. In the combined analysis, the inclusion of SR information strengthens the constraint from $0.117$ to $0.031$ eV, which corresponds to a $2\sigma$ detection of the minimum neutrino mass of $0.06$ eV.

### v2

@@ II.3 Neutrino prescription
Galaxy clusters are non-linear objects that form at the peaks of the initial density field. The Press-Schechter (PS) theory[23] firstly provides a universal formulation for describing the HMF Eq.(2): $f(\nu)=\sqrt{\frac{2}{\pi}}\nu e^{-\nu^{2}/2}$. Here, $\nu\equiv\delta_{\text{crit}}/\sigma(M,z)$ specifies the peak height, a measure of the rarity of halo, and $\delta_{\text{crit}}=1.686$ is the critical density for halo collapse, weakly dependent on the underlying cosmological model. Universality means that cosmological and redshift dependences enter the equation only through the peak height $\nu$, without affecting the functional form [23, 24, 42]. However, accurate N-body simulations have revealed deviations from universality in redshift [26, 43, 44] and in cosmological model [44].
@@
Regarding cosmological models incorporating massive neutrinos, [45, 46, 47] find that the impact of massive neutrinos can be minimized by replacing the total matter density $\rho_{m}=\rho_{cdm+b+\nu}$ in Eq.(2) and Eq.(5) with the cold dark matter + baryon density $\rho_{cdm+b}$, [48, 49, 50], and replacing the total matter power spectrum $P_{m}=P_{cdm+b+\nu}$ in Eq.(3) with the cold dark matter + baryon power spectrum $P_{cdm+b}$. The physical reason behind this is that the clustering of neutrinos around the CDM overdensity is minimal and is unlikely to impact nonlinear collapse [51, 52]. By employing this CDM prescription, [47] shows that for simulations with massive neutrinos the HMF is well reproduced by the fitting formula of Tinker et al. [26], which we will utilize in this paper.
@@
According to the peak-background split (PBS) approach [53], the bias between the halo and matter overdensities $b(M,z)$ in Eq.(17) can be derived from the HMF. Therefore, the prescription with respect to the bias for cosmologies with massive neutrinos remains the same, i.e. replacing $P_{m}$ with $P_{cdm+b}$, and $\rho_{m}$ with $\rho_{cdm+b}$, which can recover the halo bias on large scales for simulations with massive neutrinos [45, 46]. We follow [45] and continue to use the fitting formula proposed by Tinker et al. [27] for halo bias.
@@
However, this prescription only partially accounts for the effects of massive neutrinos on the large-scale bias. Massive neutrinos also introduce scale dependence in the bias due to two effects, both of which have been observed in N-body simulations: (i) halos trace CDM fluctuations rather than total (CDM + neutrino) matter fluctuations [45, 46], and (ii) massive neutrinos introduce scale-dependence in the growth of CDM density perturbations [54, 55]. The former can be eliminated by excluding neutrinos when defining bias, i.e. replacing $P_{m}$ in Eq.(II.2) with $P_{cdm+b}$. While, the latter, called the growth-induced scale-dependent bias (GISDB) [56], requires careful treatment. To incorporate the GISDB effect, we employ a fitting formula from Eq.(49) of [57], which is based on the spherical collapse model and the PBS approach [58, 57]. This method has been shown to be consistent with simulations [54]. The scale-dependence obtained by [57] is almost independent of mass and redshift [^1] The scale-dependence does show weak dependence on redshift, which is neglected in our calculation. We agree a more accurate treatment would be using Relicfast directly rather than the fitting formula of Eq. (II.3).:
@@
where $\Delta_{\Lambda CDM}=4.8\times 10^{-3}$, $\alpha=4$, $\Delta_{q}=1.6$, $\Delta_{L}=0.55f_{\nu}$ with $f_{\nu}\equiv\Omega_{\nu}/\Omega_{m}$ representing the neutrino fraction, $q=5k/k_{\text{fs}}$ (related to the free streaming scale Eq.(1)). Here, $k_{\text{eq}}$ represents the scale of matter-radiation equality [60] [^2] Here, for simplicity, we treat $k_{eq}$ as a constant., and $k_{\text{ref}}$ is a reference wave number on large scales, which we set to be $10^{-4}{{\,h^{-1}{\rm Mpc}}}$. The superscript $L$ denotes Lagrangian bias. Since we use $P_{cdm+b}$ to calculate the cluster power spectrum, the Eulerian bias is given by $b=1+b^{L}$. Finally, we can replace the scale-independent effective bias $b_{\text{eff}}(z)=b_{\text{eff}}(k_{\text{ref}},z)$ in Eq.(17) by the scale-dependent effective bias:
@@
Ignoring these two neutrino-induced scale dependences in the bias, will have an impact on cosmological parameter inference in galaxy survey [62, 63, 64, 56]. In this paper, we focus on the latter effect, i.e. GISDB effect in cluster survey.
@@ II.4 Fisher matrix
We employ the Fisher matrix techniques to assess the constraining power of clusters. It is defined as the ensemble average of the second derivatives of the logarithm of likelihood function with respect to the parameters $(p_{\mu},p_{\nu})$:
@@
where $\mathcal{L}$ represents the likelihood function. The inverse of the Fisher matrix, denoted as $\mathcal{F}^{-1}$, provides the best possible covariance matrix for the measurement errors on the parameters.
Assuming the observed cluster number count $D$ obeys a Poisson distribution with an expected value and variance $N$ [65], the logarithm of the likelihood function $\mathcal{L}(\bm{D})$ can be expressed as:
@@
where the subscripts $\{i,j\}$ represent the two-dimensional $\vec{k}^{\text{fid}}$-space cells and the observed redshift bins, respectively.
@@ II.5 Parameters
Table 1: Parameters and their fiducial values.
@@
$\Omega_{m}h^{2}$ Matter density 333including massive neutrinos 0.14240
@@
$A_{\lambda}$ Amplitude in $\ln\lambda^{ob}\right>$ 444see Eq.(II.1) 79.8
$B_{\lambda}$ Mass dependence in $\ln\lambda^{ob}\right>$ 4 0.93
$C_{\lambda}$ Redshift dependence in $\ln\lambda^{ob}\right>$ 4 -0.49
$D_{\lambda}$ Intrinsic scatter in $\sigma_{\ln\lambda^{ob}}$ 555see Eq.(13) 0.217
$\kappa$ Projection effect parameter 666see Eq.(14) 0.01
$B_{M0}$ Constant term in $M_{\text{bias}}$ 777see Eq.(11) 0
$\alpha$ Redshift dependence in $M_{\text{bias}}$ 7 0
@@
The basic cosmological model we consider is an extension of the flat $\Lambda$CDM model by adding one species of massive neutrino, labeled as $\nu\Lambda$CDM model, with the other two species assumed to be massless. We also consider $\nu w_{0}w_{a}$ model, incorporating dynamic dark energy (DE) with an equation of state (EoS) represented by $w(a)=w_{0}+(1-a)w_{a}$ [68, 69]. The set of cosmological parameters encompasses:
@@
which denotes the baryon density, matter density, neutrino density, DE density, DE Eos parameters, scalar power spectrum index and present-day normalization of the primordial power spectrum, respectively. The fiducial values for $\{\Omega_{b}h^{2},\Omega_{m}h^{2},\Omega_{DE},n_{s},\sigma_{8}\}$ follows [60], while $\Omega_{\nu}h^{2}=0.0006442$ corresponding to $M_{\nu}=0.06$ eV through [7]:
@@
Additionally, there are a set of parameters associated with the scaling relation (SR) uncertainty:
@@
where the first four parameters $\{A_{\lambda},B_{\lambda},C_{\lambda},D_{\lambda}\}$ pertain to the mass-richness scaling in Eq.(II.1) and Eq.(13), and their fiducial values follow Table IV and Eq.(B1) of [36]: { 79.8, 0.93, -0.49, 0.217}. For the projection effect term in Eq.(14), we set $\kappa=0.1^{2}$ accounting for a possible contamination rate of about 10% at low redshift for massive clusters for CSST [30, 70]. The mass bias parameters in Eq.(11) are specified as $B_{M0}=0$ and $\alpha=0$.
@@
For CSST, the solid angle is $\Delta\Omega=5.33$ sr, corresponding to the wide survey area of $17500$ deg2. The cluster redshift uncertainty in Eq.(8) is assumed to be $\sigma_{z}=0.001$ [71], considering that the photometric redshift uncertainty for galaxies is expected to be around $0.02$ for CSST [72]. The incorporation of additional slitless grating spectroscopic data will leads to a smaller error, and the redshift uncertainty of each cluster as a whole would be reduced after averaging the redshifts of the member galaxies. Thus, we adopt an optimistic value of $0.001$. The observed redshift is assumed to be up to $z_{\max}=1.5$, using either redshift-based method [73, 74] or color-based method [75]. The observed mass threshold is $M_{200m}\geq 0.836\times 10^{14}{{\,h^{-1}{\rm{M_{\odot}}}}}$, converted from $M_{500c}\geq 0.7\times 10^{14}\,\rm{M_{\odot}}$ to guarantee a completeness of $\geq 90\%$ and a purity of $\geq 90\%$ for the CSST cluster sample [73].
The fiducial wave number for cluster power spectrum spans from $k^{\text{fid}}_{\perp,\|}=0.005$ to $0.15$  Mpc-1, with a bin size of $\Delta k^{\text{fid}}_{\perp,\|}=0.005$  Mpc-1. For cluster number counts, the observed mass bin size is $\Delta\ln M^{ob}[\,\rm{M_{\odot}}]$ = 0.2, and the observed redshift bin size is $\Delta z^{ob}=0.05$. While for the cluster power spectrum, we choose $\Delta z^{ob}=0.2$, ensuring adequate number of clusters within each redshift bin.
By summing over the aforementioned bins, we obtain the Fisher matrices $\mathcal{F}_{\mu\nu}^{N}$ for the cluster number counts Eq.(27) and $\mathcal{F}_{\mu\nu}^{P}$ for power spectrum Eq.(28). Finally, we can calculate the $1\,\sigma$ marginalized error on a parameter $p_{\mu}$ as $\sigma(p_{\mu})=\sqrt{(\mathcal{F}^{-1})_{\mu\mu}}$, as shown in Table 2.
@@
Table 2: Marginalized $1\,\sigma$ errors forecasted for all the parameters described in subsection II.5. When a parameter is kept fixed, we use ’–’. The header specifies the different cosmological models ($\nu\Lambda$CDM and $\nu w_{0}w_{a}$), scenarios for our knowledge of the scaling relation (self-calibration and known SR), and probes (cluster number counts $N$, cluster power spectrum $P$, and their combination $N+P$).There are almost no constraints from $P$ on the SR parameters, denoted by ’/’, and adding priors to these parameters does not affect the main results. For complete description, please refer to the main text.
@@
$A_{\lambda}$ 1.5e+02 / 1.1e+01 - - - 2.0e+02 / 1.2e+01 - - -
$B_{\lambda}$ 1.024 / 0.203 - - - 1.356 / 0.207 - - -
$C_{\lambda}$ 1.628 / 0.328 - - - 2.294 / 0.336 - - -
$D_{\lambda}$ 0.363 / 0.074 - - - 0.415 / 0.075 - - -
$\kappa$ 0.466 / 0.129 - - - 0.539 / 0.136 - - -
$B_{M0}$ 0.860 / 0.029 - - - 1.006 / 0.033 - - -
$\alpha$ 0.431 / 0.045 - - - 0.434 / 0.047 - - -
@@
In this section, we present the main results forecasted for the CSST. The redshift distribution of the CSST clusters is shown in Figure 1, with a total number count of 408,037. (This is slightly different from our previous prediction of 415,000 in [30] due to the slight difference in the fiducial parameters.) The predicted $1\,\sigma$ marginalized errors for all parameters, especially for the total neutrino mass $M_{\nu}$, are shown in Table 2, for different probes, cosmological models, and scenarios. Specifically, these constraints are derived from the cluster number counts $N$, the cluster power spectrum $P$, and their combination $N+P$, and for the two cosmological models: the $\nu\Lambda$CDM model and the $\nu w_{0}w_{a}$ model. To address the impact of the major systematic in cluster cosmology, i.e. the uncertainty in the SR, we consider two scenarios: (1) the self-calibration scenario, where the SR parameters are jointly constrained with the cosmological parameters; and (2) the known SR scenario, where the SR parameters are fixed.
For clarity, the constraints on $M_{\nu}$ obtained from $N$, $P$ and $N+P$ are denoted as $\sigma^{N}(M_{\nu})$, $\sigma^{P}(M_{\nu})$ and $\sigma^{N+P}(M_{\nu})$ respectively.
@@ III.1 $\nu\Lambda$CDM model
Figure 2: Marginalized constraints on cosmological parameters for the $\nu\Lambda$CDM model. All contours show the 68% confidence regions. Different colors (red, blue and black) represent different probes: the cluster number counts $N$, cluster power spectrum $P$, and their combination $N+P$ in turn. Solid and dotted lines represent the self-calibration and known SR scenarios respectively.
We begin with the $\nu\Lambda$CDM model, whose constraints are shown in columns 2-7 in Table 2. The corresponding contour plot is shown in Fig. 2 for the cosmological parameters $\{M_{\nu},\Omega_{b}h^{2},\Omega_{m}h^{2},\Omega_{DE},n_{s},\sigma_{8}\}$.
It is evident that constraints on all cosmological parameters primarily come from $P$, except for $\Omega_{DE}$ in the known SR scenario. This is because $N$ is sensitive to the integration of the cdm+b power spectrum, rather than its shape, resulting in weaker constraints on $\Omega_{b}h^{2}$, $n_{s}$ and $M_{\nu}$ [40, 20, 76]. Consequently, the constraint on $M_{\nu}$ from $N$ primarily stems from the influence of neutrinos on the amplitude of the cdm+b power spectrum, leading to a strong degeneracy between $M_{\nu}$ and $\sigma_{8}$. In summary, the constraint on $M_{\nu}$ from $N$ is very weak, $\sigma^{N}(M_{\nu})=5.215\ (0.706)$ eV in the self-calibration (known SR) scenario.
The power spectrum $P$ proves to be a more sensitive probe of the neutrino mass compared to $N$, providing a constraint of $\sigma^{P}(M_{\nu})=0.201\ (0.194)$ eV on its own. When combined with $N$, the improvement factor in the self-calibration (known SR) scenario is 1.72 (6.26).
The incorporation of SR information has a substantial impact on $N$, resulting in an enhancement of constraints on all cosmological parameters by a factor of at least four. Particularly, the constraint on $M_{\nu}$ experiences a significant improvement by a factor of 7.4, shifting from $5.215$ to $0.706$ eV. In contrast, the influence of SR information on $P$ is less significant, which is reflected in the large errors of SR parameters in the third column, since the SR parameters only affect $b_{\text{eff}}$. Consequently, the improvement in the constraint on $M_{\nu}$ is only a factor of 1.04, from $0.201$ to $0.194$ eV, when incorporating SR information. In the combined analysis, the inclusion of SR information strengthens the constraint from $0.117$ to $0.031$ eV, which corresponds to a $\sim 2\sigma$ detection of the minimum neutrino mass of $0.06$ eV.